OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form 1. (-4,3) and (-2,1) 2. (2,7) and (5,2) 3. (4,2) and (2,-5) SHOW WORK

OpenStudy (campbell_st):

find the gradient of each pair...\[m = (y _{2} -y_{1})/(x _{2}-x _{1})\] then use the point slope formula \[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

1-3/-2+4 = -1............y=-x+c ( substitute the x & y of any ordered pair ( -4,3) or (-2,1) 3=-(-4)+c 3-4=c c= -1 so the equation is : y=-x-1 and do the same for the other equations